The route used by a certain motorist in commuting to work contains two distant intersections with (long) traffic signals. Let A be the event that she must stop at the first signal, and let B be the event that she must stop at the second signal.

(a) If P(A ∪ B) = 0.80, P(A) = 2P(B), and P(A ∩ B) = 0.20, what is the probability that she must stop at the first traffic signal?

(b) Suppose that P(A) = 0.30, and P(B′|A) = 0.20, what is the probability that she must stop at both traffic signals?

(c) If P (A) = 0.55, P (B) = 0.45, and P (A ∩ B) = 0.30, what is the probability that she does not have to stop when commuting to work at either traffic signal?
Answer #1

a)

P(A)+P(B) =P(A u B)+P(A n B) =0.8+0.2 =1

or P(A)+P(A)/2 =1

P(A) =2/3

b)

probability that she must stop at both traffic signals =P(A n B) =P(A)*P(B|A) =0.3*(1-0.20) =0.24

c)

probability that she does not have to stop when commuting to work at either traffic signal

=1-P(A u B) =1-(P(A)+P(B)-P(A n B)) =1-(0.55+0.45-0.30)=0.30